### v1

@@ 4 Apportionment Sequences for More Than Two Parties
The properties of apportionment sequences for two parties carry over to apportionment sequences for more than two parties because of the property of consistency. Consistency requires that the apportionment between two states is unaffected by the addition of additional states. More precisely, suppose $h$ seats are allocated among $n$ parties according to $F(h,\mathbf{p})=\mathbf{a}$ and suppose $N^{\prime}$ is a subset of parties whose total number of seats in this allocation is $\sum_{i\in N^{\prime}}a_{i}=h^{\prime}$. If we remove these seats and reapportion the remaining $h-h^{\prime}$ seats among the parties not in $N^{\prime}$, their allocation is unchanged.
@@
For any divisor method, the subsequence with $i$ and $i$ as its terms is the same as the 2-party sequence when $i$ is identified with 1 and $j$ is identified with 2. By Lemma 7, the sequence with 1s and 2s cannot have two 2s in a row. Thus, two $j$s cannot appear between consecutive $i$s. ∎
@@
Let $p_{1}=16$, $p_{2}=11$, and $p_{3}=7$ and $c\in[0,\frac{1}{9})$. Applying Proposition 11, we list the the pairwise sequential apportionment sequences for $p_{1},p_{2}$ and $c\in[0,1/5)$ (first row in Table 3); for $p_{1},p_{3}$ and $c\in[0,1/9)$ (second row in Table 3); and for $p_{2},p_{3}$ and $c\in[0,1/4)$ (third row in Table 3). We will combine the 2-party sequences to form a periodic 3-party sequence of period $h=p_{1}+p_{2}+p_{3}$ for $c\in[0,\frac{1}{9})$, the intersection of $0,\frac{1}{5})$, $0,\frac{1}{9})$, and $0,\frac{1}{4})$. Notice that rows 1, 2, and 3 of Table 3 are of length 27, 23, and 18, because each pair of $p_{i}$ and $p_{j}$ are relatively prime and the sequences are periodic of length $p_{i}+p_{j}$.
@@
We refer to the process described in Example 7 as lifting. More generally, given $n$ parties with votes $p_{i}$ for $i=1$ to $n$ and cut point $c$, we can define an algorithm to lift the $\binom{n}{2}$ 2-party apportionment sequences corresponding to $c$ to an $n$-party apportionment sequence for $c$. The algorithm begins by generating the $\binom{n}{2}$ 2-party apportionment sequences. The key step is that there is always a unique number between 1 and $n$ that begins n1n-1 of the sequences. This is the next seat to be awarded in the $n$-party apportionment sequence. After awarding this seat, that number is removed from its leftmost position in the n1n-1 sequences and the process is repeated.
@@
The following theorem proves that the algorithm generates the $n$-party apportionment sequence. In particular, the proof explains why there is always a number that appears n1n-1 times as the leftmost entry of the 2-party sequences and that the sequence terminates with all 2-party sequences’ entries being emptied.
@@
First, we show that at every step of the algorithm there are n1n-1 leftmost entries that are the same. Assume that $k$ seats have been allocated in the $n$-party apportionment sequence and that each party $i$ receives $a_{i}$ seats. The party that receives seat $k+1$ is the party $\ell$ that satisfies $\arg\max_{i\in N}\frac{p_{i}}{a_{i}-1+c}$. If more than one party satisfies the condition, then we apply the tie-break rule that smallest party number receives the seat. There are n1n-1 2-party apportionment sequences that have $\ell$ as one of the two parties. Because $\frac{p_{\ell}}{a_{\ell}-1+c}\geq\frac{p_{i}}{a_{i}-1+c}$ for all $i$, then $\ell$ appears as the leftmost entry in those n1n-1 sequences. Hence, there is always a party that is leftmost in n1n-1 sequences, as required in Algorithm 1. This party is clearly unique since in the 2-sequence between $\ell$ and any other party, $i$, party $i$ will not appear next.
Second, we show that $\frac{p_{1}+p_{2}+\cdots+p_{n}}{\gcd(p_{1},p_{2},\ldots,p_{n})}$ seats are allocated and that no extra terms are left in the 2-party sequences. Each time a seat is awarded, then the (same) leftmost characters are removed n1n-1 2-party sequences. For parties $i$ and $j$, the 2-party sequence from the algorithm consists of $\frac{p_{i}+p_{j}}{\gcd(p_{1},p_{2},\ldots,p_{n})}$ terms. Hence, there are a total of
$\sum_{i\neq j}\frac{p_{i}+p_{j}}{\gcd(p_{1},p_{2},\ldots,p_{n})}=(n-1)\cdot\frac{p_{1}+p_{2}+\cdots+p_{n}}{\gcd(p_{1},p_{2},\ldots,p_{n})}$
terms. Since each step of the algorithm removes n1n-1 terms, there are $\frac{p_{1}+p_{2}+\cdots+p_{n}}{\gcd(p_{1},p_{2},\ldots,p_{n})}$ seats awarded as required. ∎
@@
To determine all possible sequential apportionments for $p_{1}=16$, $p_{2}=11$, and $p_{3}=7$, we can apply Theorem 18 for all $c$ by comparing the pairwise sequential apportionments for $p_{1},p_{2}$ (Table 4), for $p_{1},p_{3}$ (Table 1 with the 2s replaced by 3s), and for $p_{2},p_{3}$ (Table 4). The combined apportionments appear in Table 5; this table includes colored entries to highlight how the sequences changes as the value of $c$ increases.
@@
Consistency also allows us to characterize all apportionment sequences for $n$ parties corresponding to stationary divisor methods.
@@
Given an apportionment sequence of $n$ parties $S=s_{1},s_{2},s_{3},\ldots$, let $S_{i,j}$ be the subsequence created by extracting all the $i$’s and $j$s, for all $i<j$, and let $S^{\prime}_{i,j}$ correspond to $S_{i,j}$ with $i$ and $j$ replaced by $1$ and $2$ respectively. Then $S$ corresponds to a stationary divisor method if and only if $S_{i,j}^{\prime}$ satisfies equations (1) and (2) from Proposition 8 for all $i<j$ and if the intervals (3) have nonempty intersection. If that is the case, then $S$ corresponds to the stationary divisor method for all $c$ in that intersection.
@@ 4.1 Counting the number of distinct apportionment sequences for $n$ parties
We can also determine the number of distinct apportionment sequences for $n$ parties. Recall from Section 3, if $p_{1}$ and $p_{2}$ are relatively prime with $p_{1}>p_{2}$, the interval $[0,1)$ can be partitioned into $p_{1}-p_{2}$ intervals of the form $\frac{\ell}{p_{1}-p_{2}},\frac{\ell}{p_{1}-p_{2}})$. Let $A=\{0,\frac{1}{p_{1}-p_{2}},\frac{2}{p_{1}-p_{2}},\ldots,\frac{p_{1}-p_{2}-1}{p_{1}-p_{2}}\}$ be the set of endpoints of these intervals (not including 1). Then the number of apportionment sequences is equal to $|A|+1$, where the extra $1$ corresponds to $c=1$. If $p_{1}$ and $p_{2}$ are not relatively prime, the same thing is true, where $p_{1}$ and $p_{2}$ are replaced with $p_{1}/\gcd(p_{1},p_{2})$ and $p_{2}/\gcd(p_{1},p_{2}).$
Now suppose there are $n$ parties. Each of the $\binom{n}{2}$ pairs generates its own partition of the interval $[0,1)$ and its own set of endpoints $A$. When we combine these pairs to enumerate the sequences for all $n$ parties, as in Example 8, we obtain a refinement of each of the partitions of $[0,1)$ based on the union of the sets $A$.
To determine the endpoints in this union, we require the following Lemma, which follows easily from induction.
@@
Let $A_{i}=\{0,\frac{1}{m_{i}},\ldots,\frac{m_{i}-1}{m_{i}}\}$ for some integers $m_{i}$, $i=1,\ldots,M$. Then for all $1\leq i_{1}<i_{2}<\cdots<i_{k}\leq M,$
@@
Counting the number of possible sequential apportionments for $c\in[0,1)$ is then an application of the principle of inclusion-exclusion. To make the inclusion-exclusion cleaner, let $T=\{1,2,\ldots,\binom{n}{2}\}$ be an enumeration of all possible pairs of states (the order is irrelevant). For each $k\in T$, if $k$ corresponds to $p_{i}$ and $p_{j}$ with $p_{i}>p_{j}$, then let $m_{k}=(p_{i}-p_{j})/\gcd(p_{i},p_{j})$ and let $A_{k}=\{0,1/m_{k},2/m_{k},\ldots,(m_{k}-1)/m_{k}\}$. We assume first that each of the $p_{i}$ is distinct.
@@
Let $p_{1},p_{2},\dots,p_{n}$ be the vote totals of $n$ parties where $p_{1}>p_{2}\cdots>p_{n}$. The number of possible sequential apportionments is
$1+\left|\bigcup_{k=1}^{\binom{n}{2}}A_{k}\right|$ $=1+\sum_{k=1}^{\binom{n}{2}}(-1)^{k+1}\left(\sum_{1\leq i_{1}<\cdots<i_{k}\leq\binom{n}{2}}|A_{i_{1}}\cap\cdots\cap A_{i_{k}}|\right)$
$=1+\sum_{k=1}^{\binom{n}{2}}(-1)^{k+1}\left(\sum_{1\leq i_{1}<\cdots<i_{k}\leq\binom{n}{2}}\gcd(m_{i_{1}},\ldots,m_{i_{k}})\right).$
@@
Let $\mathbf{p}=(25,17,13,5)$. Then,
$m_{1}=8\quad m_{2}=12\quad m_{3}=4\quad m_{4}=4\quad m_{5}=12\quad m_{6}=8.$
Applying Theorem 21 and not reducing fractions so it is clear where the term originates, $\left|\cup_{k=1}^{6}A_{k}\right|=48-72+80-60+24-4=16$ from the refined partition of $[0,1)$ given by $\cup_{k=1}^{6}A_{k}$ which is equal to
$\{0,1/12,1/8,2/12,1/4,4/12,3/8,5/12,2/4,7/12,5/8,8/12,3/4,10/12,7/8,11/12\}.$
Thus, there are $17$ possible sequential apportionments under stationary divisor methods.
If the party votes are not distinct, Theorem 21 can easily be modified. Note that if two parties have equal votes (say $p_{i}=p_{i+1}$), then the addition of party $i+1$ does not introduce any new endpoints to the 2-party apportionment sequences with parties with non-equal vote totals. Additionally, whenever one of the parties with equal vote totals is awarded a seat (the one with the least index by the tie-breaking rule), then the next largest indexed party with the same vote total gets the next seat, and this continues until all parties with equal vote totals get a seat. This leads to the following corollary.
@@
(to Theorem 21). Let $p_{1},p_{2},\dots,p_{n}$ be the vote totals of parties with $p_{i}\geq p_{j}$ for $i<j$. Let $N^{*}$ be one of the largest subsets of $N$ with all $p_{i}$ distinct. Relabel the elements in $N^{*}$ as $q_{1},q_{2},\ldots,q_{m}$ with $q_{i}>q_{j}$ for $i<j$. Then, then the number of possible $n$-party apportionment sequences of the parties in $N$ is the number of possible $m$-party apportionment sequences of the parties in $N^{*}$.

### v2

@@ 4 Apportionment Sequences for More Than Two Parties
The properties of apportionment sequences for two parties carry over to apportionment sequences for more than two parties because stationary divisor methods are consistent. Consistency requires that the apportionment between two parties is unaffected by the removal of another party. More precisely, suppose $h$ seats are allocated among $n$ parties according to $F(h,\mathbf{p})=\mathbf{a}$ and suppose $N^{\prime}$ is a subset of parties whose total number of seats in this allocation is $\sum_{i\in N^{\prime}}a_{i}=h^{\prime}$. If we remove these seats and reapportion the remaining $h-h^{\prime}$ seats among the parties not in $N^{\prime}$, their allocation is unchanged.
@@
For any divisor method, the subsequence with $k$ and $\ell$ as its terms is the same as the 2-party sequence when $k$ is identified with 1 and $\ell$ is identified with 2. By Lemma 1, the sequence with 1s and 2s cannot have two 2s in a row. Thus, two $\ell$’s cannot appear between consecutive $k$’s. ∎
@@
Let $p_{1}=16$, $p_{2}=11$, and $p_{3}=7$ and $c\in[0,\frac{1}{9})$. Applying Proposition 4, we list the patterns of the periodic pieces of the pairwise apportionment sequences for $p_{1},p_{2}$ and $c\in[0,\frac{1}{5})$ (first row in Table 3); for $p_{1},p_{3}$ and $c\in[0,\frac{1}{9})$ (second row in Table 3); and for $p_{2},p_{3}$ and $c\in[0,\frac{1}{4})$ (third row in Table 3). We will combine the 2-party sequences to form the periodic portion of a 3-party sequence of period $h=p_{1}+p_{2}+p_{3}=34$ for $c\in[0,\frac{1}{9})$, the intersection of $0,\frac{1}{5})$, $0,\frac{1}{9})$, and $0,\frac{1}{4})$.
@@
Notice that rows 1, 2, and 3 of Table 3 are of length $27$, $23$, and $18$, because each pair of $p_{i}$ and $p_{j}$ are relatively prime and the sequences are periodic of length $p_{i}+p_{j}$. It is not an accident that $\frac{27+23+18}{2}=34$. This ensures that all numbers of the 2-party sequences were crossed out in the construction of the 3-party sequence.
We refer to the process described in Example 8 as lifting. More generally, given $n$ parties with votes $p_{i}$ for $i=1$ to $n$ and cutpoint $c$, we can define an algorithm to lift the $\binom{n}{2}$ 2-party apportionment sequences corresponding to $c$ to an $n$-party apportionment sequence for $c$. The algorithm begins by generating the $\binom{n}{2}$ 2-party apportionment sequences. Because of the tie-breaking rule and how $p_{i}\geq p_{j}$ for $i<j$, then the n1n-1 sequences that contain $1$ (and one of the other party’s numbers between $2$ and $n$) all begin with a $1$. Once this $1$ is used to begin the sequence, then the key observation is that there is always a unique number between $1$ and $n$ that begins n1n-1 of the sequences. The party with this number is awarded the next seat in the $n$-party apportionment sequence. After awarding this seat, that number is removed from its leftmost position in the n1n-1 sequences and the process is repeated. This process is formally defined in Algorithm 1.
@@
The following theorem proves that the algorithm generates the $n$-party apportionment sequence.
@@
First, we show that at every step of the algorithm there are n1n-1 leftmost entries that are the same. Assume that $k$ seats have been allocated in the $n$-party apportionment sequence and that each party $i$ receives $a_{i}$ seats. The party that receives seat $k+1$ is a party $\ell$ that satisfies $\operatorname*{argmax}_{i\in N}\frac{p_{i}}{a_{i}+c}$. If more than one party satisfies the condition, then we apply the tie-break rule so that the party with the smallest party number receives the seat. There are n1n-1 2-party apportionment sequences that have $\ell$ as one of the two parties. Because $\frac{p_{\ell}}{a_{\ell}+c}\geq\frac{p_{i}}{a_{i}+c}$ for all $i$, then $\ell$ appears as the leftmost entry in those n1n-1 sequences. Each number appears in exactly n1n-1 sequences, so no other number can appear as the leftmost entry in n1n-1 sequences. Hence, there is always a unique party that is leftmost in n1n-1 sequences, as required in Algorithm 1.
Let $m=\gcd(p_{1},p_{2},\ldots,p_{n})$. Now we show that $\frac{p_{1}+p_{2}+\cdots+p_{n}}{m}$ seats are allocated and that no extra terms are left in the 2-party sequences. For parties $i$ and $j$, it is possible that $p_{i}$ and $p_{j}$ are not relatively prime and because of the relationship to the other $p_{k}$, it may be necessary to have more than $\frac{p_{i}+p_{j}}{\gcd(p_{i},p_{j})}$ terms in the finite sequence. To use the algorithm, we require the first $\frac{p_{i}+p_{j}}{m}$ terms from the apportionment sequence for parties $i$ and $j$. It follows that there are a total of
$\sum_{i\neq j}\frac{p_{i}+p_{j}}{m}=(n-1)\cdot\frac{p_{1}+p_{2}+\cdots+p_{n}}{m}$
terms among the 2-party apportionment sequences. Since each step of the algorithm removes n1n-1 terms, the first $\frac{p_{1}+p_{2}+\cdots+p_{n}}{m}$ seats are awarded from the period $\frac{p_{1}+p_{2}+\cdots+p_{n}}{m}$ apportionment sequence. ∎
@@
To determine all possible sequential apportionments for $p_{1}=16$, $p_{2}=11$, and $p_{3}=7$, we can apply Theorem 6 for all $c$ by comparing the pairwise sequential apportionments for $p_{1},p_{2}$ (Table 4), for $p_{1},p_{3}$ (Table 1 with the 2s replaced by 3s), and for $p_{2},p_{3}$ (Table 4). The combined apportionments appear in Table 5; this table includes colored entries to highlight how the sequences change as the value of $c$ increases.
@@
Consistency also allows us to work backwards: if given a sequence, we can apply previous 2-party results to determine if the sequence was generated by a stationary divisor method. Notice that we do not require any information about the number of votes parties receive and we are not given a possible cutpoint $c$. The following theorem acts as a characterization of all apportionment sequences for $n$ parties generated by stationary divisor methods. For simplicity, Theorem 7 excludes the case in which $q_{i}=q_{j}$. We address this case in the discussion before Corollary 5.
@@
For a periodic apportionment sequence of $n$ parties $S=s_{1},s_{2},s_{3},\ldots$ of period $Q$, let $q_{i}$ be the number of times that $i$ appears in the first $Q$ terms. Assume $q_{i}>q_{j}$ for $i<j$. Let $S_{i,j}$ be the subsequence created by extracting all the $i$’s and $j$’s, for all $i<j$, and let $S^{\prime}_{i,j}$ correspond to $S_{i,j}$ with $i$ and $j$ replaced by $1$ and $2$, respectively.
Then $S$ is generated by a stationary divisor method if and only if $S_{i,j}^{\prime}$ for all $i<j$ satisfies Equations (1) and (2) from Proposition 5 with $p_{i}=q_{i}/\gcd(q_{i},q_{j})$ and $p_{j}=q_{j}/\gcd(q_{i},q_{j})$. If that is the case, then $S$ is generated by a stationary divisor method with cutpoint $c$ for all $c$ in that intersection of the intervals in Equation (3).
@@ 4.1 Counting the number of distinct apportionment sequences for $n$ parties
In this subsection, we determine the number of distinct apportionment sequences for $n$ parties. Recall from Section 3, if $p_{1}$ and $p_{2}$ are relatively prime with $p_{1}>p_{2}$, the interval $[0,1)$ can be partitioned into $p_{1}-p_{2}$ intervals of the form $\frac{\ell}{p_{1}-p_{2}},\frac{\ell}{p_{1}-p_{2}})$. Let $A=\{0,\frac{1}{p_{1}-p_{2}},\frac{2}{p_{1}-p_{2}},\ldots,\frac{p_{1}-p_{2}-1}{p_{1}-p_{2}}\}$ be the set of left endpoints of these intervals. Then the number of apportionment sequences is equal to $|A|+1$, where the extra $1$ corresponds to $c=1$. If $p_{1}$ and $p_{2}$ are not relatively prime, the same statement holds but with $p_{1}$ and $p_{2}$ replaced by $p_{1}/\gcd(p_{1},p_{2})$ and $p_{2}/\gcd(p_{1},p_{2})$, respectively.
Now suppose there are $n$ parties where all the $p_{i}$ are distinct; we address the case where the $p_{i}$ are not distinct separately. Each of the $\binom{n}{2}$ pairs generates its own partition of the interval $[0,1)$ and its own set of left endpoints. When we combine these pairs to enumerate the sequences for all $n$ parties, as in Example 9, we obtain a refinement of each of the partitions of $[0,1)$ based on the union of the sets of leftmost endpoints from all the pairs.
Let $M=\{1,2,\ldots,\binom{n}{2}\}$ be the set of indices of all possible pairs of parties (the order is irrelevant). For each $k\in M$, if $k$ corresponds to $p_{i}$ and $p_{j}$ with $p_{i}>p_{j}$, then let $m_{k}=(p_{i}-p_{j})/\gcd(p_{i},p_{j})$ and let $A_{k}=\{0,1/m_{k},2/m_{k},\ldots,(m_{k}-1)/m_{k}\}$. It follows that the number of possible sequential apportionments under all stationary divisors is
$\left|\bigcup_{k\in M}A_{k}\right|+1,$
where the “+1+1” comes from $c=1$. To count the number of elements in the union, we apply the principle of inclusion/exclusion in Theorem 8, which will keep track of double, triple, etc. counting, which happens when the $m_{k}$ are not distinct. The following lemma gives conditions for an $m_{k}$ to appear in every set of a collection of $A_{\ell}$.
@@
Let $\mathbf{p}=(p_{1},\ldots,p_{n})$ with $p_{i}\neq p_{j}$ for all $i$ and $j$. Let $A_{k}=\{0,\frac{1}{m_{k}},\ldots,\frac{m_{k}-1}{m_{k}}\}$ for some $k\in M$, where $m_{k}=\frac{p_{r}+p_{s}}{\gcd(p_{r},p_{s})}$ for some $r$ and $s$. Then, for all $1\leq i_{1}<i_{2}<\cdots<i_{k}\leq\binom{n}{2}$,
@@
To see that the lemma holds, consider the two cases: the $m_{i_{j}}$ share a common factor $m>1$ and the $m_{i_{j}}$ are relatively prime collectively. If $m=\gcd(m_{i_{1}},\ldots,m_{i_{k}})>1$, then the leftmost endpoint $\frac{1}{m}$ appears in each $A_{i_{j}}$, as well as every other $\frac{\ell}{m}$ for $\ell=2$ to m1m-1. If the $m_{i_{j}}$ are relatively prime collectively, then $A_{i_{1}}\cap\cdots\cap A_{i_{k}}=\{0\}$, and the sets only share the leftmost endpoint 0. The following theorem requires Lemma 2 and is a direct application of the principle of inclusion/exclusion [4, p.165].
@@
Let $p_{1},p_{2},\dots,p_{n}$ be the vote totals of $n$ parties where $p_{1}>p_{2}>\cdots>p_{n}$. The number of possible sequential apportionments from stationary divisor methods is
$1+\left|\bigcup_{k=1}^{\binom{n}{2}}A_{k}\right|$ $=1+\sum_{\{i_{1},\ldots,i_{k}\}\subset M}(-1)^{k+1}|A_{i_{1}}\cap\cdots\cap A_{i_{k}}|$
$=1+\sum_{\{i_{1},\ldots,i_{k}\}\subset M}(-1)^{k+1}\gcd(m_{i_{1}},\ldots,m_{i_{k}}),$
where $gcd(m_{i_{j}})$ is defined to be $m_{i_{j}}$.
The following example demonstrates the application of Theorem 8, as well as shows all intervals of cutpoints that give rise to distinct apportionment sequences.
@@
Let $\mathbf{p}=(25,17,13,5)$. Then, $M=\{1,2,3,4,5,6=\binom{4}{2}$}. Define
$m_{1}=\textstyle\frac{25-17}{1}=8\quad$ $m_{2}=\textstyle\frac{25-13}{1}=12\quad$ $m_{3}=\textstyle\frac{25-5}{5}=4$
$m_{4}=\textstyle\frac{17-13}{1}=4\quad$ $m_{5}=\textstyle\frac{17-5}{1}=12\quad$ $m_{6}=\textstyle\frac{13-5}{1}=8.$
Because $m_{1}=m_{6}$, $m_{2}=m_{5}$, and $m_{3}=m_{4}$, then $\cup_{k=1}^{6}A_{k}=\cup_{k=1}^{3}A_{k}$. To give insight into the proof of Theorem 8, we show how the principle of inclusion/exclusion works for $\cup_{k=1}^{3}A_{k}$:
$|A_{1}\cup A_{2}\cup A_{3}|=|\{0,\textstyle\frac{1}{8},\textstyle\frac{2}{8},\ldots,\textstyle\frac{7}{8}\}\cup\{0,\textstyle\frac{1}{12},\textstyle\frac{2}{12},\ldots,\textstyle\frac{11}{12}\}\cup\{0,\textstyle\frac{1}{4},\textstyle\frac{2}{4},\textstyle\frac{3}{4}\}|$
$=\left(|A_{1}|+|A_{2}|+|A_{3}|\right)-\left(|A_{1}\cap A_{2}|+|A_{1}\cap A_{3}|+|A_{2}\cap A_{3}|\right)-|A_{1}\cap A_{2}\cap A_{3}|$
$=\left(8+12+4\right)-\left(4+4+4\right)+4=16.$
This value is verifiable because we can count the number of elements in $\cup_{k=1}^{3}A_{k}$ directly:
$|\cup_{k=1}^{3}A_{k}|=|\textstyle{\left\{0,\frac{1}{12},\frac{1}{8},\frac{2}{12},\frac{1}{4},\frac{4}{12},\frac{3}{8},\frac{5}{12},\frac{2}{4},\frac{7}{12},\frac{5}{8},\frac{8}{12},\frac{3}{4},\frac{10}{12},\frac{7}{8},\frac{11}{12}\right\}}|=16.$
Alternatively, we can apply Theorem 8 to show $\left|\cup_{k=1}^{6}A_{k}\right|=48-72+80-60+24-4=16$, as expected. Thus, there are $16+1=17$ possible sequential apportionments under stationary divisor methods, when the D’Hondt apportionment for $c=1$ is included. Each of the following intervals for $c$ give a distinct apportionment sequence:
$\textstyle{\left[0,\frac{1}{12}\right),\left[\frac{1}{12},\frac{1}{8}\right),\left[\frac{1}{8},\frac{2}{12}\right),\left[\frac{2}{12},\frac{1}{4}\right),\left[\frac{1}{4},\frac{4}{12}\right),\left[\frac{4}{12},\frac{3}{8}\right),\left[\frac{3}{8},\frac{5}{12}\right),\left[\frac{5}{12},\frac{2}{4}\right),}$
$\textstyle{\left[\frac{2}{4},\frac{7}{12}\right),\left[\frac{7}{12},\frac{5}{8}\right),\left[\frac{5}{8},\frac{8}{12}\right],\left[\frac{8}{12},\frac{3}{4}\right),\left[\frac{3}{4},\frac{10}{12}\right),\left[\frac{10}{12},\frac{7}{8}\right),\left[\frac{7}{8},\frac{11}{12}\right),\{1\}.}$
If the party votes are not distinct, Theorems 8 and 7, as well as other results, can easily be modified. Note that if two parties have equal votes (say $p_{i}=p_{i+1}$), then the addition of party $i+1$ does not introduce any new endpoints to the 2-party apportionment sequences with parties with non-equal vote totals. Additionally, whenever one of the parties with equal vote totals is awarded a seat (the one with the least index by the tie-breaking rule), then the next largest indexed party with the same vote total gets the next seat, and this continues until all parties with equal vote totals get a seat. This leads to the following corollary.
@@
Let $p_{1},p_{2},\dots,p_{n}$ be the vote totals of parties with $p_{i}\geq p_{j}$ for $i<j$. Let $N^{*}$ be a largest subset of $N$ with all $p_{i}$ distinct. Relabel the elements in $N^{*}$ as $q_{1},q_{2},\ldots,q_{m}$ with $q_{i}>q_{j}$ for $i<j$. Then, the number of possible $n$-party apportionment sequences of the parties in $N$ is the number of possible $m$-party apportionment sequences of the parties in $N^{*}$.